# Functional bottlenecks can emerge from non-epistatic underlying traits

### v1

@@ Abstract
Protein fitness landscapes frequently exhibit epistasis, where the effect of a mutation depends on the genetic context in which it occurs, i.e., the rest of the protein sequence. Epistasis increases landscape complexity, often resulting in multiple fitness peaks. In its simplest form, known as global epistasis, fitness is modeled as a non-linear function of an underlying additive trait. In contrast, more complex epistasis arises from a network of (pairwise or many-body) interactions between residues, which cannot be removed by a single non-linear transformation. Recent studies have explored how global and network epistasis contribute to the emergence of functional bottlenecks - fitness landscape topologies where two broad high-fitness basins, representing distinct phenotypes, are separated by a bottleneck that can only be crossed via one or a few mutational paths. Here, we introduce and analyze a simple model of global epistasis with an additive underlying trait. We demonstrate that functional bottlenecks arise with high probability if the model is properly calibrated. Our results underscore the necessity of sufficient heterogeneity in the mutational effects selected by evolution for the emergence of functional bottlenecks. Moreover, we show that the model agrees with experimental findings, at least in small enough combinatorial mutational spaces.
@@ I Introduction
What is sometimes called ‘global’ epistasis is based on introducing the simplest possible non-linearity in the genotype-phenotype-fitness mapping De Visser and Krug (2014); Tenaillon (2014); Pahujani and Krug (2025); Hwang et al. (2017); Otwinowski (2018); Otwinowski et al. (2018); Poelwijk et al. (2019); Husain and Murugan (2020); Reddy and Desai (2021); Phillips et al. (2021); Johnson et al. (2023); Buda et al. (2023). More precisely, one assumes the existence of an underlying additive phenotype (or ‘trait’) $E({\bf a})=\sum_{i=1}^{L}h_{i}(a_{i})$ associated with each genotype ${\bf a}$. Under this assumption, the variation of phenotype $E({\bf a})$ associated with substituting $a$ with $b$ at site $i$, $\Delta E_{i}(a\rightarrow b)=h_{i}(b)-h_{i}(a)$, is independent of the rest of the genotype. Next, one assumes that the fitness $F({\bf a})=\phi(E({\bf a}))$ is a non-linear function of $E({\bf a})$, for instance a sigmoid (Fig. 1a). Under these assumptions, while the variation of the underlying trait is independent of the rest of the sequence, the variation of fitness depends on it due to the non-linearity of $\phi$. Yet, such non-linearity is relatively easy to handle, because one can deduce the function $\phi(E)$ by ‘deconvolution’ of experimental fitness measurements Otwinowski (2018); Otwinowski et al. (2018); Poelwijk et al. (2019); Phillips et al. (2021); Buda et al. (2023), and thus describe the fitness function with a limited number of parameters. In the example above, two parameters enter in the definition of $\phi(E)$, and of the order of $L$ parameters enter in the definition of $E$. Furthermore, such a globally epistatic fitness landscape only features a single maximum, because $F=\phi(E)$ is an increasing function of $E$, and $E$ itself has a single maximum, obtained by choosing at each site the value $a_{i}$ that maximizes the additive contribution $h_{i}(a_{i})$ Pahujani and Krug (2025). Multi-peaked landscapes can be obtained by choosing a non-monotonic function $\phi(E)$, as in Fisher’s Geometric Model De Visser and Krug (2014); Tenaillon (2014); Hwang et al. (2017); Pahujani and Krug (2025).
@@
A more complex behavior is obtained in presence of what is sometimes called ‘network’ epistasis Lunzer et al. (2010); Olson et al. (2014); Figliuzzi et al. (2016); Rivoire et al. (2016); Poelwijk et al. (2016); Cocco et al. (2018); Starr et al. (2018); Ballal et al. (2020); Lyons et al. (2020); Miton et al. (2021); Bakerlee et al. (2022); Vigué et al. (2022); Chen et al. (2024), which implies that no function $E=\phi^{-1}(F)$ can transform fitness into a purely additive function of the genotype Poelwijk et al. (2019); Phillips et al. (2021); Buda et al. (2023). While a function $\phi$ can be chosen to make $E$ as close as possible to an additive form, higher-order interactions will still play a role in the presence of network epistasis. More formally, $E$ can always be decomposed as a sum of additive, two-site, three-site, and higher-order interactions Weinreich et al. (2013); Poelwijk et al. (2016); Sailer and Harms (2017b, a); Domingo et al. (2018); Poelwijk et al. (2019); Ballal et al. (2020); Phillips et al. (2021); Buda et al. (2023), i.e.,
$E({\bf a})=\sum_{i}h_{i}(a_{i})+\sum_{i,j}J_{ij}(a_{i},a_{j})+\sum_{i,j,k}K_{ijk}(a_{i},a_{j},a_{k})+\cdots$ (1)
@@
however, due to network epistasis, no choice of $\phi$ can eliminate the higher-order terms $J_{ij}$, $K_{ijk}$, and beyond. Let us focus for simplicity on the case in which only the two-site interactions are present. The number of parameters required to fully specify the function $E$ now scales as the number of pairs, i.e., proportionally to $L^{2}$, which makes inferring the fitness function from a few measurements significantly more difficult. Even more importantly, a function like (1) can have many local maxima, corresponding to many fitness peaks, leading to a rough, complex fitness landscape.
@@
The relative importance of global and network epistasis has been a subject of study in recent years. A particular focus of the debate has been the role played by epistasis in shaping the topology of the fitness landscape in the proximity of functional switches. Poelwijk et al. Poelwijk et al. (2019) studied experimentally the fitness landscape separating two phenotypically distinct variants of the Entacmaea quadricolor fluorescent protein, one fluorescent in the red and the other in the blue, separated by 13 mutations. Assaying all $2^{13}$ intermediate variants, they observed, in addition to global epistasis, significant network contributions (both pairwise and higher order) to fitness. They also identified a functional bottleneck separating the ‘blue’ and ‘red’ space of fluorescent variants – a narrow region of accessible evolutionary paths between the two phenotypes – which they attributed to epistatic constraints. These functional bottlenecks illustrate how fitness landscapes can remain navigable despite their ruggedness. While epistasis can constrain viable evolutionary paths, it can also facilitate adaptation by carving out narrow high-fitness pathways that connect otherwise isolated functional states – typically via critical “switch” mutations Weinreich et al. (2006); Wu et al. (2016); Kantroo et al. (2024).
@@
In a more recent study, Alqatari and Nagel Alqatari and Nagel (2024) used similar methods to investigate function switch in an elastic network, considered as a model for allosteric behavior in proteins. Analyzing ensembles of fitness landscapes, they observed different topologies at the threshold for viable evolution, and found that functional bottlenecks – often involving a single critical mutation – emerged generically, similar to the observations in the experimental study.
@@
Here, we identify the minimal conditions under which such topological features emerge. Motivated by these works, we investigate whether network epistasis is a necessary condition for the existence of functional bottlenecks in fitness landscapes. We thus introduce a simple toy model of global epistasis, based on a random underlying additive trait and a non-linear fitness function. We investigate the random fitness landscapes generated by our toy model after proper calibration, and show that there is a wide range of parameters such that the resulting random topologies have a high probability to display a functional bottleneck. This result, by itself, shows that global epistasis alone is sufficient to generate functional bottlenecks. Furthermore, we studied the role of the distribution of the single mutational effects (SMEs). More specifically, we find that when the model is properly calibrated, this distribution of SMEs selected during evolution displays strong heterogeneity.
Our model thus sheds light on some fundamental constraints that shape fitness landscapes and possibly evolution. Finally, the model is analytically solvable to some extent, which might help making further progress in the investigation of its behavior for larger mutational spaces and more than two phenotypic variants.
@@ II Results
II.1 Reanalyzing experimental data
To obtain inspiration for our toy model, we begin by reanalyzing experimental data from Poelwijk et al. Poelwijk et al. (2019). In this work, the whole space of intermediates between two variants of the Entacmea Quadricolor fluorescent protein eqFP611 were studied. The two reference wild types are mTagBFP2, which is fluorescent in the red, and mKate2 which is fluorescent in the blue. Their sequences differ by 13 amino acids, resulting in a total of $81922^{13}=$ 8192 different combinations of intermediate amino acid substitutions, whose fluorescence was measured in the experiment.
@@ II.1 Reanalyzing experimental data
Figure 1: Reanalyzing experimental data from Ref. Poelwijk et al. (2019). (a) Schematic shape of the fitness functions for the red and blue phenotypes as a function of the underlying trait $E$. Sigmoid functions have been used for illustrations. The two wild type structures with the corresponding mutations are also shown, from Ref. Poelwijk et al. (2019). (b) Histogram of the value of $E$ obtained from Eqs. (2,3) for each of the $2^{13}$ variants. The red and blue lines correspond respectively to the wild type values $E^{WT}_{R}$ and $E^{WT}_{B}$. (c) Histogram of the absolute value $|\Delta E|$ of single mutational effects (SMEs) from (5) for all $13\times 2^{13}$ single mutations that can be obtained from the dataset, shown in log-log (main panel) and lin-log (inset) scales. The black curve is the fit obtained with the Pareto distribution in (6). (d) Topology of the space of paths obtained keeping only genotypes ${\bf a}$ with $|E({\bf a})|>E_{C}$, and finding the largest possible value of $E_{C}=0.53$ (dashed line) such that the red and blue wild types remain connected. The lower panel reports the values of $E({\bf a})$ for each functional genotype (blue dots for $E>E_{C}$ and red dots for $E<-E_{C}$) as a function of the number of mutations $m$ from the blue wild type, with the gray lines connecting pairs of genotypes that differ by a single mutation. The upper panel shows the resulting graph of connections.
@@
We downloaded the original data from Ref. Poelwijk et al. (2019) and represented the genotypes with binary vectors of 13 variables, with ${\bf a}_{B}=(0,0,\cdots,0)$ being the blue mKate2 and ${\bf a}_{R}=(1,1,\cdots,1)$ being the red mTagBFP2. In the experiment, E.Coli cells expressing the mutant genotypes were encapsulated into droplets and sorted by a microfluidic device able to push droplets into separate channels according to their fluorescence. From the raw experimental measurement, we assign to each genotype a red fitness, $F_{R}({\bf a})$, and a blue fitness, $F_{B}({\bf a})$, based on the enrichment in the two channels, each normalized to its corresponding wild type value such that $F_{R}({\bf a}_{R})=F_{B}({\bf a}_{B})=1$ (see the SI for details). This normalization provides a consistent reference scale, with each wild type serving as the baseline functional state for its phenotype, enabling a more direct quantitative comparison between the red and blue fitness values. Note that here, fitness refers to fluorescence intensity, which is unrelated to reproductive success. Rather, fluorescence measures how well the protein performs its function. Under the assumption that better-functioning proteins offer a selective advantage in their relevant environments, fluorescence serves as a reasonable proxy for fitness.
@@
For each measured phenotype (we will use a suffix ‘R’ for red and ‘B’ for blue fluorescence), we can define an underlying trait and a non-linear mapping to fitness ${F=\phi(E)}$. Inverting the non-linear function, we can thus derive the underlying phenotype from the measured fitness,
$E_{R}({\bf a})=\phi_{R}^{-1}\left[F_{R}({\bf a})\right]\ ,\quad E_{B}({\bf a})=\phi_{B}^{-1}\left[F_{B}({\bf a})\right]\ ,$ (2)
@@
where we used for simplicity the same non-linear function $\phi_{R}^{-1}(x)=\phi_{B}^{-1}(x)=x^{0.44}$ that was chosen in Ref. Poelwijk et al. (2019) to minimize the epistatic contributions. We note that when there are two phenotypes only, for simplicity, we can follow Alqatari and Nagel Alqatari and Nagel (2024) and encode them both in a single trait
@@
which is positive for the ‘blue’ phenotype and negative for the ‘red’ phenotype. This is only possible when the two phenotypes are exclusive, as in Ref. Poelwijk et al. (2019), which may not be true in other cases Bendixsen et al. (2019); Phillips et al. (2021). We also define the wild type phenotype values as
$E^{WT}_{B}=E({\bf a}_{B})\ ,\qquad E^{WT}_{R}=E({\bf a}_{R})\ .$ (4)
The structure of the two proteins and the 13 substitutions are illustrated in Fig. 1a, together with a schematic illustration of the non-linear mapping between $E=E_{B}-E_{R}$ and the two fitnesses. Note that other fitting functions could be used in the analysis, such as the sigmoid function $\phi$ (used for illustration in Fig. 1a), as suggested by more recent studies of global epistasis Otwinowski et al. (2018); Phillips et al. (2021); Buda et al. (2023). However, it would be even better to infer the non-linearity directly from the fitting procedure using the rank statistics Carlson et al. (2025).
Fig. 1b shows the histogram of $E({\bf a})$ across all $2^{13}$ genotypes, with red and blue lines corresponding to the two wild types. Note that the wild type values are $E^{WT}_{B}\sim 0.88$ and $E^{WT}_{R}\sim-0.89$. Their absolute value slightly differs from one due to the contribution from the other phenotype. As observed in Ref. Poelwijk et al. (2019), most genotypes are non-functional with $E\sim 0$. There is a rather sharp peak of blue genotypes around $E\sim 1$, while red genotypes exhibit a much broader distribution, ranging from $E\sim-0.5$ down to large negative values of $E$. This indicates that some intermediates are significantly more fluorescent in the red than the red wild type itself.
@@
$p(x)=\frac{\alpha}{2x_{m}(\alpha+1)}\times\begin{cases}1&\text{if }|x|<x_{m}\ ,\\ \big{(}\frac{x_{m}}{|x|}\big{)}^{\alpha+1}&\text{if }|x|>x_{m}\ ,\\ \end{cases}$ (6)
@@
Finally, in Fig. 1d, we examine the topology of the resulting functional bottleneck. To do so, we first need a definition of functional genotype. As observed in Ref. Poelwijk et al. (2019) and evident from Fig. 1b, the threshold for defining a genotype as ‘functional’ is somewhat arbitrary, and playing with this threshold changes the topology substantially Alqatari and Nagel (2024). In this work, we consider that the fitness function is growing fast around a given threshold $E_{\rm th}$ (either on the positive or negative side, see Fig. 1a). Hence, genotypes with $|E({\bf a})|<E_{\rm th}$ are considered dysfunctional and evolutionarily disallowed. Because the wild types are functional, they must of course be located above the functionality threshold, i.e. $|E^{WT}|\gtrsim E_{\rm th}$. However, Fig. 1b shows that no path of single mutations exists that connects the two wild types while always maintaining $|E({\bf a})|\geq E^{WT}$, hence we need to consider $E_{\rm th}<E^{WT}$ in order to preserve the connection between the wild types. (Note that a different result was obtained in Ref. Poelwijk et al. (2019) when the two phenotypes were not normalized to the wild types. We discuss this point in the SI.) What is, then, the largest value of the functionality threshold $E_{\rm th}$ such that at least one viable mutational path exists between the two wild types?

### v2

@@ Abstract
Protein fitness landscapes frequently exhibit epistasis, where the effect of a mutation depends on the genetic context in which it occurs, i.e., the rest of the protein sequence. Epistasis increases landscape complexity, often resulting in multiple fitness peaks. In its simplest form, known as global epistasis, fitness is modeled as a non-linear function of an underlying additive trait. In contrast, more complex epistasis arises from a network of (pairwise or many-body) interactions between residues, which cannot be removed by a single non-linear transformation. Recent studies have explored how global and network epistasis contribute to the emergence of functional bottlenecks - fitness landscape topologies where two broad high-fitness basins, representing distinct phenotypes, are separated by a bottleneck that can only be crossed via one or a few mutational paths. Here, we introduce and analyze a stylized model of global epistasis with an additive underlying trait. We demonstrate that functional bottlenecks arise with high probability if the model is properly calibrated. Furthermore, our results underscore that a proper balance between neutral and non-neutral mutations is needed for the emergence of functional bottlenecks.
Author Summary
A central challenge in the study of protein evolution is understanding how interactions between mutations influence evolutionary dynamics. These interactions, collectively known as epistasis, play a key role in shaping protein ’fitness landscapes’, representing maps between amino acid sequences and their functional performance. However, the impact of epistasis on landscape ruggedness and accessibility remains a subject of debate.
Recent experiments have revealed the existence of functional bottlenecks: regions of low fitness that restrict evolutionary transitions between proteins with different functionalities. While these bottlenecks represent a significant constraint on evolutionary outcomes and are often attributed to complex networks of interacting mutations, we demonstrate that they can arise in a much simpler setting.
Using a stylized model where fitness depends nonlinearly on an underlying additive trait, we show that functional bottlenecks do not require complex network epistasis. Instead, they can emerge from the inherent variability of the mutational effect distribution. Specifically, these bottlenecks arise when the mutational landscape is dominated by small, nearly neutral effects, but remains punctuated by enough strongly non-neutral mutations to create sharp fitness transitions. Our findings offer a new perspective on the constraints shaping protein evolution and highlight the critical role of mutational effect heterogeneity in determining evolutionary accessibility.
@@ I Introduction
What is sometimes called ‘global’ epistasis is based on introducing the simplest possible non-linearity in the genotype-phenotype-fitness mapping De Visser and Krug (2014); Tenaillon (2014); Pahujani and Krug (2025); Hwang et al. (2017); Otwinowski (2018); Otwinowski et al. (2018); Poelwijk et al. (2019); Husain and Murugan (2020); Reddy and Desai (2021); Phillips et al. (2021); Johnson et al. (2023); Buda et al. (2023). More precisely, one assumes the existence of an underlying additive phenotype (or ‘trait’) $E({\bf a})=\sum_{i=1}^{L}h_{i}(a_{i})$ associated with each genotype $\bf a$}. Under this assumption, the variation of phenotype $E({\bf a})$ associated with substituting $a$ with $b$ at site $i$, $\Delta E_{i}(a\rightarrow b)=h_{i}(b)-h_{i}(a)$, is independent of the rest of the genotype. Next, one assumes that the fitness $F({\bf a})=\phi(E({\bf a}))$ is a non-linear function of $E({\bf a})$, for instance a sigmoid (Fig. 1a). Under these assumptions, while the variation of the underlying trait is independent of the rest of the sequence, the variation of fitness depends on it due to the non-linearity of $\phi$. Yet, such non-linearity is relatively easy to handle, because one can deduce the function $\phi(E)$ by ‘deconvolution’ of experimental fitness measurements Otwinowski (2018); Otwinowski et al. (2018); Poelwijk et al. (2019); Phillips et al. (2021); Buda et al. (2023); Park et al. (2024); Dupic et al. (2024); Carlson et al. (2025), and thus describe the fitness function with a limited number of parameters. In concrete examples, few parameters enter in the definition of $\phi(E)$, and of the order of $L$ parameters enter in the definition of $E$. Furthermore, such a globally epistatic fitness landscape only features a single maximum, because $F=\phi(E)$ is an increasing function of $E$, and $E$ itself has a single maximum, obtained by choosing at each site the value $a_{i}$ that maximizes the additive contribution $h_{i}(a_{i})$ Pahujani and Krug (2025). Multi-peaked landscapes can be obtained by choosing a non-monotonic function $\phi(E)$, as in Fisher’s Geometric Model De Visser and Krug (2014); Tenaillon (2014); Martin and Lenormand (2015); Hwang et al. (2017); Pahujani and Krug (2025).
@@
A more complex behavior is obtained in presence of what is sometimes called ‘network’ epistasis Lunzer et al. (2010); Olson et al. (2014); Figliuzzi et al. (2016); Rivoire et al. (2016); Poelwijk et al. (2016); Cocco et al. (2018); Starr et al. (2018); Ballal et al. (2020); Russ et al. (2020); Lyons et al. (2020); Miton et al. (2021); Bakerlee et al. (2022); Vigué et al. (2022); Chen et al. (2024); Park et al. (2024); Dupic et al. (2024); Carlson et al. (2025), which implies that no function $E=\phi^{-1}(F)$ can transform fitness into a purely additive function of the genotype Poelwijk et al. (2019); Phillips et al. (2021); Buda et al. (2023). While a function $\phi$ can be chosen to make $E$ as close as possible to an additive form, higher-order interactions will still play a role in the presence of network epistasis. More formally, $E$ can always be decomposed as a sum of additive, two-site, three-site, and higher-order interactions Weinreich et al. (2013); Poelwijk et al. (2016); Sailer and Harms (2017b, a); Domingo et al. (2018); Poelwijk et al. (2019); Ballal et al. (2020); Phillips et al. (2021); Buda et al. (2023), i.e.,
$\begin{split}E({\bf a})=&\sum_{i}h_{i}(a_{i})+\sum_{i,j}J_{ij}(a_{i},a_{j})\\ &+\sum_{i,j,k}K_{ijk}(a_{i},a_{j},a_{k})+\cdots\end{split}$ (1)
@@
however, due to network epistasis, no choice of $\phi$ can eliminate the higher-order terms $J_{ij}$, $K_{ijk}$, and beyond. Let us focus for simplicity on the case in which only the two-site interactions are present. The number of parameters required to fully specify the function $E$ now scales as the number of pairs, i.e., proportionally to $L^{2}$, which makes inferring the fitness function from a few measurements significantly more difficult. Even more importantly, the function in Eq. (1) can have many local maxima, corresponding to many fitness peaks, leading to a rough, complex fitness landscape. From Eq. (1), any nonlinear function of $E({\bf a})$ – that is, any form of global epistasis – can, in principle, be represented as a (possibly infinite) sum of high-order terms on the right-hand side. However, this representation is less practical, as it requires estimating a large number of parameters. A more effective strategy is to first capture as much of the nonlinearity as possible using a single global function. This substantially reduces the residual nonlinearity, making it easier to model the remainder with only low-order local terms.
@@
The relative importance of global and network epistasis, and how to properly fit $E({\bf a})$, has recently been the subject of intense debate Park et al. (2024); Dupic et al. (2024); Carlson et al. (2025). A particular focus of the debate has been the role played by epistasis in shaping the topology of the fitness landscape in the proximity of functional switches. For instance, Poelwijk et al. Poelwijk et al. (2019) studied experimentally the fitness landscape separating two phenotypically distinct variants of the Entacmaea quadricolor fluorescent protein, one fluorescent in the red and the other in the blue, separated by 13 mutations. Assaying all $2^{13}$ intermediate variants, they observed, in addition to global epistasis, significant network contributions (both pairwise and higher order) to fitness. They also identified a functional bottleneck separating the ‘blue’ and ‘red’ space of fluorescent variants, i.e. a narrow region of accessible evolutionary paths between the two phenotypes, which they attributed to epistatic constraints. These functional bottlenecks illustrate how fitness landscapes can remain navigable despite their ruggedness. While epistasis can constrain viable evolutionary paths, it can also facilitate adaptation by carving out narrow high-fitness pathways that connect otherwise isolated functional states – typically via critical “switch” mutations Weinreich et al. (2006); Wu et al. (2016); Kantroo et al. (2024).
@@
In a subsequent study, Alqatari and Nagel Alqatari and Nagel (2026) used similar methods to investigate functional switches in an elastic network, considered as a model for allosteric behavior in proteins. Analyzing ensembles of fitness landscapes, they observed different topologies at the threshold for viable evolution, and found that functional bottlenecks – often involving a single critical mutation – emerged generically, similar to the observations in the experimental study.
Many other recent studies investigated small combinatorial fitness landscapes. However, in most cases, these landscapes connect a reference wild type to an evolved variant with increased or decreased fitness with respect to a single functionality Buda et al. (2023); Papkou et al. (2023); Johnston et al. (2024). Experimental studies that investigate the combinatorial landscape separating two reference genotypes with distinct functionality are less widespread Westmann et al. (2024); Starr et al. (2017); Bendixsen et al. (2019); Phillips et al. (2021). Moreover, in many of these studies functional transitions proceed through promiscuous intermediate variants that are as fit, or even fitter, than the two reference genotypes. Allowing for functional promiscuity represents a less constrained scenario than one where a functional switch must occur between mutually exclusive phenotypes. In the latter case, which is the one we consider here, the requirement for specialization naturally makes the landscape more prone to the formation of bottlenecks. The question of how widespread bottlenecks are – particularly in presence of promiscuity – is not addressed here.
@@
In this work, we aim to identify the minimal conditions under which bottleneck topologies can emerge. Specifically, we investigate whether network epistasis is a necessary condition for the existence of functional bottlenecks in fitness landscapes. To address this, we introduce a simple, stylized model of global epistasis based on a random underlying additive trait and a nonlinear fitness function. Our model is not intended to fit specific empirical data; rather, it follows the spirit of the Fisher’s Geometric Model and related frameworks De Visser and Krug (2014); Tenaillon (2014); Martin and Lenormand (2015); Hwang et al. (2017); Pahujani and Krug (2025) by generating an ensemble of random fitness landscapes with stylized features. Upon calibration, we demonstrate that our model encompasses a wide range of parameters for which the resulting topologies exhibit a functional bottleneck with high probability. This result, by itself, shows that global epistasis alone is sufficient to generate functional bottlenecks.
Furthermore, we studied the role of the distribution of single mutational effects (SMEs). More specifically, we find that when the model is properly calibrated, the distribution of SMEs selected by evolution displays a proper balance between nearly neutral SMEs and enough strongly non-neutral SMEs, which is needed to create sharp fitness transitions. Our stylized model thus sheds light on some fundamental constraints that shape fitness landscapes and possibly evolution.
@@ II Results
The central result of our work is the introduction of a stylized fitness landscape model, designed to investigate the minimal conditions under which bottleneck topologies emerge. However, to provide the motivation and inspiration for this construction, we begin with a pedagogical discussion of experimental data from Poelwijk et al. Poelwijk et al. (2019). This analysis of empirical data serves to ground the stylized features of the model that follows.
II.1 Experimental data on fitness bottlenecks
Poelwijk et al. Poelwijk et al. (2019) analyzed a family of variants of the naturally occurring wild type Entacmea Quadricolor red fluorescent protein eqFP611 Wiedenmann et al. (2005). More specifically, from a natural variant of eqFP611, called eqFP578 (with 76% amino acid sequence homology), the TagRFP mutant was engineered via random mutagenesis. TagRFP is 21 mutations away from eqFP578 (Merzlyak et al., 2007, see SI) and has enhanced fluorescence and stability. From TagRFP, two engineered variants have been derived to have distinct fluorescence [^1] See https://www.fpbase.org/protein/eqfp578 for the complete tree of variants.:
  • mTagBFP2, which is fluorescent in the blue and has 13 mutations and 6 insertions relative to TagRFP Subach et al. (2011);
  • mKate2, which is fluorescent in the deep red and has 8 mutations and one insertion relative to TagRFP Shcherbo et al. (2009).
The sequences of mTagBFP2 and of mKate2 differ by 13 mutations (12 of which are shown in Fig. 1a reproduced from Ref. Poelwijk et al. (2019)), thus resulting in a total of $2^{13}=8192$ different combinations of intermediate amino acid substitutions. The experiment of Poelwijk et al. Poelwijk et al. (2019) measured the fluorescence of all these intermediates, both in the red and in the blue. Note that there are also 5 insertions in mTagBFP2 with respect to mKate2 that were neglected in the experiment (details were not found in Ref. Poelwijk et al. (2019)).
We focus on these published data as they characterize a combinatorial landscape bridging two reference genotypes with distinct functionalities, within which a functional bottleneck was explicitly identified. A more recent study, employing a similar experimental design but investigating a different protein, reported the absence of such a bottleneck Westmann et al. (2024). This discrepancy highlights the topological variability of fitness landscapes and motivates our search for the minimal conditions that govern the emergence of these constraints.
@@ II.1 Experimental data on fitness bottlenecks
Figure 1: Experimental data from Ref. Poelwijk et al. (2019). (A) Schematic shape of the fitness functions for the red and blue phenotypes as a function of the underlying trait $E$. Sigmoid functions have been used for illustrations. The two reference structures with the corresponding mutations are also shown, from Ref. Poelwijk et al. (2019). (B) Histogram of the value of $E$ obtained from Eqs. (2,3) for each of the $2^{13}$ variants. The red and blue lines correspond respectively to the reference values $E^{\rm{ref}}_{R}$ and $E^{\rm{ref}}_{B}$. (C) Histogram of the absolute value $|\Delta E|$ of single mutational effects (SMEs) from Eq. (5) for all $13\times 2^{13}$ single mutations that can be obtained from the dataset, shown in log-log (main panel) and lin-log (inset) scales. The black curve is the fit obtained with the Pareto distribution in Eq. (6). (D) Topology of the space of paths obtained keeping only genotypes $\bf a$} with $|E({\bf a})|>E_{C}$, and finding the largest possible value of $E_{C}=0.53$ (dashed line) such that the red and blue reference sequences remain connected. The lower panel reports the values of $E({\bf a})$ for each functional genotype (blue dots for $E>E_{C}$ and red dots for $E<-E_{C}$) as a function of the number of mutations $m$ from the blue reference, with the gray lines connecting pairs of genotypes that differ by a single mutation. The upper panel shows the resulting graph of connections.
@@
We downloaded the original data from Ref. Poelwijk et al. (2019) and represented the genotypes with binary vectors of 13 variables, with ${\bf a}_{B}=(0,0,\cdots,0)$ being the blue mKate2 and ${\bf a}_{R}=(1,1,\cdots,1)$ being the red mTagBFP2. In the experiment, E.Coli cells expressing the mutant genotypes were sorted by a microfluidic device able to push them into separate channels according to their fluorescence. From the raw experimental measurement, we assign to each genotype a red fitness, $F_{R}({\bf a})$, and a blue fitness, $F_{B}({\bf a})$, based on the enrichment in the two channels, each normalized to its corresponding reference value such that $F_{R}({\bf a}_{R})=F_{B}({\bf a}_{B})=1$ (see the SI for details). This normalization provides a consistent reference scale, with each reference variant serving as the baseline functional state for its phenotype, enabling a more direct quantitative comparison between the red and blue fitness values.
Note that a specific measurable trait – fluorescence intensity – is here employed as a proxy for fitness, despite its lack of direct coupling to reproductive success. Instead, fluorescence quantifies the protein’s functional efficiency. Under the assumption that enhanced biophysical performance may confer a selective advantage within the relevant environment, fluorescence serves as a proxy for fitness. This approach is standard in many empirical studies of fitness landscapes, such as Refs. Poelwijk et al. (2019); Westmann et al. (2024).
@@
For each measured phenotype (we will use a suffix ‘R’ for red and ‘B’ for blue fluorescence), we can define an underlying trait and a non-linear mapping to fitness $F=\phi(E)$}. Inverting the non-linear function, we can thus derive the underlying trait from the measured fitness,
$E_{R}({\bf a})=\phi_{R}^{-1}\left[F_{R}({\bf a})\right]\ ,\quad E_{B}({\bf a})=\phi_{B}^{-1}\left[F_{B}({\bf a})\right]\ .$ (2)
The optimal methodology for inferring the function $\phi(E)$ remains a subject of active debate Otwinowski et al. (2018); Phillips et al. (2021); Buda et al. (2023); Park et al. (2024); Dupic et al. (2024); Carlson et al. (2025); for instance, it has been suggested that the nonlinearity be inferred directly through rank-based statistics Carlson et al. (2025). While acknowledging that results can be sensitive to the specific choice of $\phi$, for our illustrative purposes we keep the same power-law form $\phi^{-1}(x)=x^{0.44}$ utilized in Ref. Poelwijk et al. (2019) for both phenotypes. Alternative specifications, such as the sigmoid function shown in Fig. 1a, may yield quantitatively different results, but the qualitative topology of the landscape remains robust.
@@
We note that when there are two phenotypes only, for simplicity, we can follow Alqatari and Nagel Alqatari and Nagel (2026) and encode them both in a single trait
@@
which is positive for the ‘blue’ phenotype and negative for the ‘red’ phenotype. This is only possible when the two phenotypes are exclusive, as in Ref. Poelwijk et al. (2019), which may not be true in other cases Bendixsen et al. (2019); Phillips et al. (2021); Westmann et al. (2024); Starr et al. (2017). We also define the reference trait values as
$E^{\rm{ref}}_{B}=E({\bf a}_{B})\ ,\qquad E^{\rm{ref}}_{R}=E({\bf a}_{R})\ .$ (4)
The structure of the two proteins and the 13 substitutions are illustrated in Fig. 1a, together with a schematic illustration of the non-linear mapping between $E=E_{B}-E_{R}$ and the two fitnesses.
Fig. 1b shows the histogram of $E({\bf a})$ across all $2^{13}$ genotypes, with red and blue lines corresponding to the two reference variants. Note that the reference values are $E^{\rm{ref}}_{B}\sim 0.88$ and $E^{\rm{ref}}_{R}\sim-0.89$. Their absolute value slightly differs from one due to the contribution from the other phenotype. As observed in Ref. Poelwijk et al. (2019), most genotypes are non-functional with $E\sim 0$. There is a rather sharp peak of blue genotypes around $E\sim 1$, while red genotypes exhibit a much broader distribution, ranging from $E\sim-0.5$ down to large negative values of $E$. This indicates that some intermediates are significantly more fluorescent in the red than the red reference variant itself.
@@
[…]
@@
Finally, in Fig. 1d, we examine the topology of the resulting functional bottleneck. To perform this analysis, we need to first establish a formal definition of a ‘functional’ genotype. As noted in Ref. Poelwijk et al. (2019) and illustrated in Fig. 1b, the threshold for functionality is inherently context-dependent; shifting this threshold can substantially alter the perceived landscape topology Alqatari and Nagel (2026). Indeed, viability is not an intrinsic property of a genotype alone but is contingent upon the environment – specifically the selection pressure exerted on the relevant trait – which varies across different experimental conditions. In this work, we consider that the fitness function is growing fast around a given threshold $E_{\rm th}$ (either on the positive or negative side, see Fig. 1a). Hence, genotypes with $|E({\bf a})|<E_{\rm th}$ are considered dysfunctional and evolutionarily disallowed. Because the reference variants are functional, they must of course be located above the functionality threshold, i.e. $|E^{\rm{ref}}|\gtrsim E_{\rm th}$. However, Fig. 1d shows that no path of single mutations exists that connects the two reference variants while always maintaining $|E({\bf a})|\geq E^{\rm{ref}}$, hence we need to consider $E_{\rm th}<E^{\rm{ref}}$ in order to preserve the connection between the reference sequences. (Note that a different result was obtained in Ref. Poelwijk et al. (2019) when the two phenotypes were not normalized to the reference values. We discuss this point in the SI.) What is, then, the largest value of the functionality threshold $E_{\rm th}$ such that at least one viable mutational path exists between the two reference variants?